A certain math club makes 30 bars of chocolate a week and sells these at P30 each. Before the chocolate can all be sold, the pupil found out that 6 bars eaten by mice. How much will be the total sale at the end of the four week month? solution
1
Expert's answer
2020-09-29T16:00:07-0400

Amount that would be made from sale of 30 bars in one week selling each bar at P30 would be, (number of bars x cost of each bar)

"(30x30) = 30^2"


In 4 weeks would be,

"(30x30)4 = 30^2 x 4=3600"


Cost of the 6 bars eaten by mice will be,

"(30x6)=180"


Therefore amount made from the sale in 4 weeks will be,

"(3600-180)= 3420"


Therefore the sale will amount = P3420